If a person is traveling at 40mph and leaves at 8 am, but on the return he/she travels at a rate of 25mph and returns at 2:30pm, how far has the person travelled?
Answer by Alan3354(69443)

If a person is traveling at 40mph and leaves at 8 am, but on the return he/she travels at a rate of 25mph and returns at 2:30pm, how far has the person travelled?
------------
1430 - 0800 = 6.5 hours
Assuming the person was moving all the time,
for a round-trip, the avg speed is
--
2*40*25/(40 + 25) = 2000/65 = 400/13 mi/hr
---
d = r*t = (400/13)*6.5
d = 200 miles